A GARDEN PLOT IS 16 FEET WIDE. IF THE PERIMETER OF THE PLOT IS 72 FEET, WHAT IS THE LENGTH OF THE TILE?

Answer by checkley79(3341)

P=2L+2W
W=16
72=2L+2*16
72=2L+32
2L=72-32
2L=40
L=40/2
L=20 FT.
PROOF:
72=2*20+2*16
72=40+32
72=72